# Large Language Models for Generative Information Extraction: A Survey

### v2

@@ 6 Evaluation & Analysis
Despite the great success of LLMs in various natural language processing tasks, their performance in the field of information extraction still have room for improvement [67]. To alleviate this problem, recent research has explored the capabilities of LLMs with respect to the major subtasks of IE, i.e., NER [217, 147], RE [194, 228], and EE [53]. Considering the superior reasoning capabilities of LLMs, Xie et al. [217] proposed four reasoning strategies for NER, which are designed to simulate ChatGPT’s potential on zero-shot NER. Wadhwa et al. [194] explored the use of LLMs for RE and found that few-shot prompting with GPT-3 achieves near SOTA performance, while Flan-T5 can be improved with chain-of-thought style explanations generated via GPT-3. For EE task, Gao et al. [53] showed that ChatGPT still struggles with it due to the need for complex instructions and a lack of robustness.
@@
7 Future Directions
8 Benchmarks & Backbones
Benchmarks As shown in Table 6, we compiled a comprehensive collection of benchmarks covering various domains and tasks, to provide researchers with a valuable resource that they can query and reference as needed. Moreover, we also summarized the download links for each dataset in our open source repository.
Backbones We briefly describe some backbones that are commonly used in the field of generative information extraction, which is shown in Table 7.

### v3

@@ 6 Evaluation & Analysis
Despite the great success of LLMs in various natural language processing tasks [236, 237], their performance in the field of information extraction still have room for improvement [193]. To alleviate this problem, recent research has explored the capabilities of LLMs with respect to the major subtasks of IE, i.e., NER [190, 83], RE [169, 170], and EE [191]. Considering the superior reasoning capabilities of LLMs, Xie et al. [190] proposed four reasoning strategies for NER, which are designed to simulate ChatGPT’s potential on zero-shot NER. Wadhwa et al. [169] explored the use of LLMs for RE and found that few-shot prompting with GPT-3 achieves near SOTA performance, while Flan-T5 can be improved with chain-of-thought style explanations generated via GPT-3. For EE task, Gao et al. [191] showed that ChatGPT still struggles with it due to the need for complex instructions and a lack of robustness.
@@
7 Future Directions
8 Benchmarks & Backbones
8.1 Representative Datasets.
Table 6: A summary of some representative IE datasets.
Dataset Summary
CoNLL03 [239] Dataset scope: NER; 1,393 English news articles from Reuters; 909 German news articles; 4 annotated entity types.
CoNLL04 [240] Dataset scope: RE; entity-relation triples from news sentences; 4 entity types; 5 relation types.
ACE05 [241] Dataset scope: NER, RE and EE; various text types and genres; 7 entity types; 7 relation types; 33 event types and 22 argument roles.
In this section, we introduce representative datasets of NER, RE and EE respectively, and show brief summary of each dataset in the Table 6 to help readers better understand these tasks.
CoNLL03. The CoNLL03 [239] is a representative dataset for NER, including 1,393 news articles in English and 909 news articles in German. The English portion of the corpus was sourced from the Shared Task dataset curated by Reuters. This dataset encompasses annotations for four distinct entity types: PER (person), LOC (location), ORG (organization), and MISC (including all other types of entities).
CoNLL04. The CoNLL04 [240] is a well-known benchmark dataset for RE tasks, comprising sentences extracted from news articles that each contain at least one entity-relation triple. It has four kinds of entities (PER, ORG, LOC, OTH) and five kinds of relations (Kill, Work For, Live In, OrgBased In, Located In).
ACE05. Automatic Content Extraction 05 [241] is widely recognized and utilized for IE tasks. It serves as a valuable resource to assess the efficacy of automated systems in extracting structured information from diverse textual sources, encompassing news articles, interviews, reports, etc. Moreover, this dataset covers a broad range of genres including politics, economics, sports, among others. Specifically for the EE task within ACE05, it comprises 599 news documents that encapsulate 33 distinct event types and 22 argument roles.
@@ 8 Benchmarks & Backbones
8.2 Benchmarks.
[…]
8.3 Backbones.
We briefly describe some backbones that are commonly used in the field of generative information extraction, which is shown in Table 8.
@@ 9 Conclusion
{acknowledgement}
This work was supported in part by the grants from National Natural Science Foundation of China (No.62222213, 62072423). Additionally, this research was partially supported by Research Impact Fund (No.R1015-23), APRC - CityU New Research Initiatives (No.9610565, Start-up Grant for New Faculty of CityU), CityU - HKIDS Early Career Research Grant (No.9360163), Hong Kong ITC Innovation and Technology Fund Midstream Research Programme for Universities Project (No.ITS/034/22MS), Hong Kong Environmental and Conservation Fund (No. 88/2022), and SIRG - CityU Strategic Interdisciplinary Research Grant (No.7020046), Huawei (Huawei Innovation Research Program), Tencent (CCF-Tencent Open Fund, Tencent Rhino-Bird Focused Research Program), Ant Group (CCF-Ant Research Fund, Ant Group Research Fund), Alibaba (CCF-Alimama Tech Kangaroo Fund (No. 2024002)), CCF-BaiChuan-Ebtech Foundation Model Fund, and Kuaishou.